please solve -3 (2x -1)= 27 - 2x Please show work. Thank you in advance.
Answer by ptth176(50)

-3 (2x -1)= 27 - 2x
=>(-3)2x-3(-1)=27-2x
=>-6x+3=27-2x
=>-4x=24
=>x=-24/4=6